OpenStudy (anonymous):

Just need some homework help please. Find the polar for of the following complex number: sqrt3 - sqrt3i. I believe the answer is 3(cos pi/4 + i sin pi/4), am I correct?

OpenStudy (anonymous):



OpenStudy (anonymous):

angle should be either \(-\frac{\pi}{4}\) or \(\frac{7\pi}{4}\)

OpenStudy (anonymous):

absolute value is \[\sqrt{\sqrt{3}^2+\sqrt{3}^2}\] \[=\sqrt{3+3}=\sqrt{6}\]

OpenStudy (anonymous):

the other answer options I have are 6(cos 7pi/4 + i sin 7pi/4) sqrt6(cos 7pi/4 - i sin 7pi/4) and sqrt6(cos 7pi/4 + i sin 7pi/4) I thought I worked it out correctly, guess not.

OpenStudy (anonymous):

it is the last one

OpenStudy (anonymous):

it helps to know what quadrant you are in so you can find the angle more easily \(\frac{\pi}{4}\) would put you in quadrant 1 but you are in quadrant 4

OpenStudy (anonymous):

Yes, after seeing what you wrote I guess so. Thank you for that.

OpenStudy (anonymous):

yw oh and don't forget \(|a+bi|=\sqrt{a^2+b^2}\)